A car is accelerating in a straight line. It's initial speed is 5 m/s and final speed is 12 m/s. The time is 4 seconds. Find the acceleration.
1
Expert's answer
2020-10-25T18:29:52-0400

By the definition of the acceleration, we have:


"a=\dfrac{v-v_0}{t}=\dfrac{12\ \dfrac{m}{s}-5\ \dfrac{m}{s}}{4\ s}=1.75\ \dfrac{m}{s^2}."

Answer:

"a=1.75\ \dfrac{m}{s^2}."